### v1

# Floquet Engineering of Relativistic Electrons using Propagating Waves

Abstract
This study examines the electronic states of quantum materials influenced by waves $V(Qx-\omega t)$ propagating through them. These waves can include laser light, polaritons, acoustic waves, and sliding density waves, whose speed may vary widely. Since these waves are periodic in both space and time, the electrons in the materials form Floquet-Bloch states when influenced by the waves. We investigate band engineering by these waves, using three-dimensional (3D) Dirac electrons in circularly polarized laser wave as a case study. We classify the fields by their speed into time, light and space-like waves, and find that the electronic states are distinctive in the three cases. We demonstrate phase transitions of the electronic bands and discover emergent Floquet Weyl bands, which potentially become Type-II in the vicinity of the transition. The effect of Floquet band engineering is enhanced by Lorentz contraction, which is maximized when the wave speed approaches the Fermi velocity. Additionally, we discuss the geometric nature of the response currents.
Introduction.— Floquet engineering, i.e., a method to coherently control system using carefully designed time-periodic driving fields, has been applied to various systems to realize exotic states of matter Eckardt (2017); Oka and Kitamura (2019); Rudner and Lindner (2020); de la Torre et al. (2021); Ozawa et al. (2019). This technique enables the study of Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011), anomalous Floquet topological insulators Rudner et al. (2013), and Floquet Weyl semimetals Wang et al. (2014); Ebihara et al. (2016); Chan et al. (2016); Hübener et al. (2017); Yoshikawa et al. (2022); Bucciantini et al. (2017); Hirai et al. (2024) as well as other exotic many-body states Harper et al. (2020); Mori (2023). While many applications assume spatially homogeneous driving fields, Floquet theory can also be applied to spatially varying drives. A propagating wave $V(Qx-\omega t)$ moving in the $x$-direction with speed $v=\omega/Q$ exemplifies this. There are numerous instances of propagating waves in quantum materials where their interaction with electrons presents intriguing challenges.These include laser light (mostly treated as homogeneous fields Wang et al. (2013); McIver et al. (2020); Ito et al. (2023)), slow light using photonic crystals Baba (2008), sliding density waves Grüner (1988); Lee et al. (1973); Monçeau et al. (1976); Bardeen (1979); Cox et al. (2008), surface acoustic waves Puebla et al. (2020), and polaritons Basov et al. (2021); Walicki et al. (2024a). The coherence and high intensity of these waves make the nonperturbative response, explored through Floquet theory Yang et al. (2024); Walicki et al. (2024b), particularly interesting.
The nonlinear response of quantum materials to external fields, especially considering geometric effects, has become a vibrant research area. Studies employing Floquet theory M. Torres and A. Kunold (2005); Oka and Aoki (2009, 2011), perturbation theory Sodemann and Fu (2015); Sipe and Shkrebtii (2000); Morimoto and Nagaosa (2016); Ahn and Nagaosa (2021); Ahn et al. (2022), and tunneling theory Takayoshi et al. (2021a); Kitamura et al. (2020) have been conducted. Generalizing the linear response theory of electric conductivity to the non-perturbative domain through propagating waves offers another pathway. In linear response theory Kubo (1957), we consider the current ($i,j=x,y,z$)
$J_{i}(\bm{r},t)=\sum_{j}\sigma_{ij}(\bm{Q},\omega)E_{j}(\bm{r},t),$ (1)
induced by an applied ac-electric field $E_{i}(\bm{r},t)=E_{i}(\bm{Q},\omega)e^{i(\bm{Q}\cdot\bm{r}-\omega t)}$. The zero-temperature conductivity $\sigma_{ij}(\bm{Q},\omega)$ is obtained by evolving the ground state |$\ket{\psi}$ under the field $V(t)$ turned on in the infinite past:
[…]
Here, $V(t)$ represents the ac-electric field. The Kubo formula, obtained by evaluating $\bra{\psi^{\prime}}j_{i}\ket{\psi^{\prime}}$ assuming adiabatic connection of |$\ket{\psi^{\prime}}$ to |$\ket{\psi}$, is expressed as a perturbative series in $V$. Several questions arise: (a) When is the perturbative picture of |$\ket{\psi^{\prime}}$ applicable? (b) In the calculation of the dc-conductivity, we take the limit $\bm{Q}\to\bm{0}$ and $\omega\to 0$ in this order. What happens if we consider the reverse order? (c) What are the new geometric properties that emerge in the nonperturbative regime? Applying Floquet theory, we can non-perturbatively study the properties of the state in an propagating wave Eqn. (2) and answer these questions. To proceed with a concrete example, this work focuses on three-dimensional (3D) Dirac electrons in propagating waves. Dirac electrons are relativisitic and its Lorentz covariance enables us to categorize the propagating waves into time, space and light-like ones according to its speed relative to the Fermi velocity. The resulting Floquet-Bloch states have distinctive properties in each of these cases, leading to unique electronic behaviors and potential phase transitions.
Floquet-Bloch theorem for electrons in propagating waves.— Let us study non-interacting three dimensional (3D) Dirac electrons  Yan and Felser (2017); Gorbar et al. (2021); Fedotov et al. (2023) in a propagating wave as a prototypical example. We employ the four vector notation of relativity theory, which is useful in non-relativistic system as well. The space-time and momentum-energy variables are expressed as $x^{\mu}=(ct,\bm{x})$ and $k^{\mu}=(\varepsilon/c,\bm{k})$, and their product is $k\cdot x=k^{\mu}x_{\mu}=\varepsilon t-\bm{k}\cdot\bm{x}$. We use Einstein’s summation convention and $v_{\mu}=\eta_{\mu\nu}v^{\nu}$ with metric $\eta_{\mu\nu}=\mbox{diag}(1,-1,-1,-1)$. The speed of light $c$ is introduced as the Fermi velocity of the Dirac electron. We consider a Lagrangian (see e.g., Ref. Bjorken and Drell (1964))
$\mathcal{L}=\bar{\psi}\left[i\gamma^{\mu}(\partial_{\mu}+ieA_{\mu}+ieA_{5\mu}\gamma^{5})-m-\gamma^{0}V(x)\right]\psi$ (3)
with fermions $\psi=^{t}(\psi^{1},\psi^{2},\psi^{3},\psi^{4}),\;\bar{\psi}=^{t}(\psi^{1},\psi^{2},\psi^{3},\psi^{4})^{\dagger}\gamma^{0}$. The $4\times 4$ gamma matrices are denoted as $\gamma^{\mu}$ and $\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$ is the chirality operator. Here, $A_{\mu}$ and $A_{5\mu}$ denotes the gauge and chiral gauge fields, and $m$ is the mass. A wave that propagates in the $x$-direction with momentum $Q$, frequency $\omega$ and speed $v=\omega/Q$ can be expressed by
$V(x)=V_{1}\cos(Qx-\omega t-\varphi_{A})+V_{2}\sin(Qx-\omega t-\varphi_{B})$ (4)
using Hermitian matrices $V_{i}\;(i=1,2)$. Generically, there can be phase variables $\varphi_{i}(x)$ known as phasons, but we ignore them in the present work and set $\varphi_{i}=0$. With the four vector notation, we can rewrite the propagating wave as
$V(x)=Ve^{-iq\cdot x}+V^{\dagger}e^{iq\cdot x}$ (5)
[…]
$H(\bm{k})=\bm{\Gamma}\cdot(\bm{k}+e\bm{A}+e\bm{A}^{5}\gamma^{5})+m\gamma^{0}-eA_{0}-eA^{5}{}_{0}\gamma^{5},$ (6)
where $\Gamma^{i}=\gamma^{0}\gamma^{i}$. We can introduce a Floquet-Bloch decomposition of the wave function as
|$\ket{\Psi(x)}=\frac{1}{\sqrt{N}}e^{-ik\cdot x}\ket{\Phi(q\cdot x)}=\frac{1}{\sqrt{N}}e^{-ik\cdot x}\sum_{n=\infty}^{\infty}\ket{\phi^{n}}e^{-inq\cdot x},$ (7)
where the Floquet-Bloch state has a periodicity of $q\cdot x\sim q\cdot x+2\pi$ ($N$ is a normalization factor). Note that this periodicity includes the spatial $x\sim x+2\pi/Q$ and temporal $T\sim T+2\pi/\omega$ periodicities when $Q\neq 0$ and $\omega\neq 0$ respectively. For the Floquet-Bloch state (Eqn. (7)) to be a solution of the time dependent Schrödinger equation, the Floquet quasi-momentum-energy vector $k^{\mu}=(\varepsilon/c,\bm{k})$ and the state |$\ket{\Phi}$ must satisfy the Floquet-Bloch equation. We employ Sambe’s space-time picture denoting the Floquet-Bloch state as an infinite dimensional vector with Fourier modes as components |$\ket{\Phi_{\bm{k}}}=^{t}(\ldots,\ket{\phi^{n-1}_{\bm{k}}},\ket{\phi^{n}_{\bm{k}}},\ket{\phi^{n+1}_{\bm{k}}},\ldots)$. The Schrödinger equation can be expressed as an eigenvalue problem that determines the $\alpha$-th eigenstate Yang et al. (2024),
$\mathcal{H}_{\bm{k}}\ket{\Phi_{\alpha\bm{k}}}=\varepsilon_{\alpha\bm{k}}\ket{\Phi_{\alpha\bm{k}}},$ (8)
$\mathcal{H}_{\bm{k}}^{nn^{\prime}}=(H_{\bm{k}+n\bm{q}}-ncq_{0})\delta_{nn^{\prime}}+V^{\dagger}\delta_{n-1n^{\prime}}+V\delta_{n+1n^{\prime}}.$ (9)
This equation can be considered as an extension of the standard Floquet equation, with the $ncq_{0}$-term, and Bloch equation, with the $+n\bm{q}$ term, mixing them into a single defining equation of the Floquet-Bloch state.
The geometric properties of this state can be captured by the Berry connection four vector $\mathcal{A}_{\alpha\bm{k}}^{\mu}=(c\mathcal{A}_{\alpha\bm{k}}^{0},\mathcal{A}_{\alpha\bm{k}}^{i})$ of the $\alpha$-th state, whose components are defined by
$\mathcal{A}_{\alpha\bm{k}}^{\mu}(q\cdot x)=\bra{\Phi_{\alpha\bm{k}}(q\cdot x)}\eta^{\mu\nu}(i\partial_{k^{\nu}})\ket{\Phi_{\alpha\bm{k}}(q\cdot x)}.$ (10)
The Berry curvature of the Floquet-Bloch state can be define as Xiao et al. (2010)
$\Omega^{\mu\nu}_{\alpha\bm{k}}(q\cdot x)=\partial_{k_{\mu}}\mathcal{A}_{\alpha\bm{k}}^{\nu}(q\cdot x)-\partial_{k_{\nu}}\mathcal{A}_{\alpha\bm{k}}^{\mu}(q\cdot x),$ (11)
which is gauge invariant. Following the derivation given in Ref. Oka and Aoki (2009) we can formulate the Floquet TKNN formula as
$\sigma_{xy}=e^{2}\int\frac{d\bm{k}}{(2\pi)^{d}}\sum_{\alpha}f_{\alpha}(\bm{k})\int_{0}^{2\pi}\frac{ds}{2\pi}\Omega^{xy}_{\alpha\bm{k}}(s)$ (12)
where we changed variable to $s=q\cdot x$ and set $d=3$. The distribution function $f_{\alpha}(\bm{k})$ that specifies the occupation of the $\alpha$-th Floquet state should be determined by an independent method such as the Floquet Keldysh Green’s function method Schüler et al. (2020); Oka and Aoki (2009), master equation Dehghani et al. (2014, 2015) or Floquet-Boltzmann equation Genske and Rosch (2015); Seetharam et al. (2019).
There are other gauge invariants of the Berry connection than Eqn. (12). The Floquet state |$\ket{\Phi_{\alpha\bm{k}}(q\cdot x)}$ has two periodicities, namely $\bm{k}\sim\bm{k}+\bm{Q}$ and $t\sim t+2\pi/\omega$, assuming $Q\neq 0$ and $\omega\neq 0$ respectively. By integrating the Berry four vector along the periodic direction, we can define the non-adiabatic polarization and geometric phase by
[…]
$\gamma_{\alpha\bm{k}}=\int_{0}^{2\pi}\frac{ds}{2\pi}\bra{\Phi_{\alpha\bm{k}}(s)}(i\partial_{s})\ket{\Phi_{\alpha\bm{k}}(s)}.$ (14)
Here, $\bm{k}_{\perp}=(0,k_{y},k_{z})$ is the crystal momentum perpendicular to the wave propagation direction $\bm{Q}$. The geometric phase is well studied in Floquet systems and are related to topological pumping (see for e.g., Ref. Nakagawa et al. (2020)). The polarization is known to be an important geometric quantity in crystals Resta (1992); King-Smith and Vanderbilt (1993) and Eqn. (13) gives a direct extension to Floquet-Bloch states in propagating waves. We note that the definition of the polarization requires that the state $\alpha$ is non-degenerate during the path $k_{x}$ from 00 to $Q$. We can define the total polarization induced by the propagating wave by
$P^{x}_{\rm tot}(Qx-\omega t)=e\int\frac{d\bm{k}_{\perp}}{(2\pi)^{d-1}}\sum_{\alpha}f_{\alpha}(\bm{k})P^{x}_{\alpha\bm{k}_{\perp}}(Qx-\omega t).$ (15)
This result shows that the longitudinal current $J^{x}_{\rm tot}(Qx-\omega t)=\frac{d}{dt}P^{x}_{\rm tot}(Qx-\omega t)$ that flows parallel to the propagating wave $V(x)$ is non-dissipative and is a geometric quantity. This current can be considered as one of the non-perturbative generalization of the Kubo formula Eqn. (1), in particular, the quantum Hall current Thouless et al. (1982); Kohmoto (1985). This can be noticed if we consider an electric field in the $y$-direction $E_{y}(Q,\omega)$ as our driving field by setting $V=\frac{E_{y}(Q,\omega)}{\omega}\Gamma^{y}$. Then, the current $J^{x}_{\rm tot}(Qx-\omega t)$ is a non-perturbative Hall current that flows in the direction perpendicular to the electric field.
Application to relativistic fermions.— Below, we focus our discussion to applications on Dirac electrons. Although we can directly study the Floquet-Bloch states in waves with general $q^{\mu}$-vector by solving Eqn. (8), we can also use the Lorentz covariance of the Dirac electron to study the band engineering systematically. To this end, we perform a coordinate transformation to a “rest frame” in which the wave is stopped.
Figure 1: (a) Using the Lorentz transformation Eqn. (22), we can move to the “rest frame” where we have a static (homogeneous) wave for the space-like (time-like) propagating waves. The solid lines ($\omega^{2}/v_{F}^{2}-Q^{2}=\mbox{const.}$.) describe the $q^{\mu}$-vectors that are related by Lorentz transformation. (b) In the lab frame, we have a wave $V$ propagating with speed $v$ and static Dirac electrons. (c) In the rest frame, the space like wave (time like wave) is stopped (homogeneous but oscillating), and Dirac electrons move in the opposite direction with speed $v$. The periodicity of the wave is enlongated by the Lorentz contraction factor $\gamma$.
We set the “speed of light” parameter used in the above formulation to be the Fermi velocity of the Dirac electron, i.e., $c=v_{F}$. In contrast to elementary particle physics, in condensed matter systems we can consider waves in materials faster than the Dirac electrons. Let us call waves that propagate faster than the Fermi velocity as time-like ($v/v_{F}>1$), equally fast as light-like ($v/v_{F}=1$), and those that are slower as space-like ($v/v_{F}<1$). We expect enhancement of relativistic effects when the speed of the propagating wave $v$ approaches $v_{F}$. We introduce a Lorentz boost $x^{\prime}=ax$ (or $x^{\prime\mu}=a^{\mu}{}_{\nu}x^{\nu}$) given by
$\left(\begin{array}[]{c}ct^{\prime}\\ x^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}\cosh\theta&-\sinh\theta\\ -\sinh\theta&\cosh\theta\end{array}\right)\left(\begin{array}[]{c}ct\\ x\end{array}\right)$ (22)
with $y^{\prime}=y,\;z^{\prime}=z$. By selecting the boost parameter $\theta$ to be $\gamma=\cosh\theta=1/\sqrt{1-v_{F}^{2}/v^{2}}$ and $\gamma=\cosh\theta=1/\sqrt{1-v^{2}/v_{F}^{2}}$ for the time and space-like waves, respectively, we can transform the frequency-wavelength parameter to
[…]
Let us call the $x^{\prime}$-frame determined this way as the rest frame (although in the time-like case the wave is actually oscillating in time), and the original $x$-frame as the lab frame (Fig. 1). The factor $\gamma$ in the above relations represents Lorentz contraction, i.e., a moving object’s length is measured to be shorter than its length measured in the rest frame. $\gamma$ will diverge as $v$ approach $v_{F}$.
Next, we relate the electron wave function in the two frames using Lorentz transformation. To this end, we enforce the fields to be related by Lorentz transformation
$A^{\prime}{}^{\mu}=a^{\mu}{}_{\nu}A^{\nu},\;A_{5}^{\prime}{}^{\mu}=a^{\mu}{}_{\nu}A_{5}{}^{\nu},\;F^{\prime\mu\nu}=a^{\mu}{}_{\rho}a^{\nu}{}_{\sigma}F^{\rho\sigma}.$ (25)
The propagating wave is also transformed according to its representation of the Lorentz group. The matrix $\gamma^{0}V$ that specifies the propagating wave can be expressed by a linear superposition of the following sixteen $4\times 4$ matrices Bjorken and Drell (1964); Peskin and Schroeder (1995),
$1,\;\gamma^{\mu},\;\sigma^{\mu\nu}=\frac{1}{2}[\gamma^{\mu},\gamma^{\nu}],\;\gamma^{\mu}\gamma^{5},\;\gamma^{5}.$ (26)
These quantities transform as a scalar, vector, spin 2 field, pseudo-vector, and a pseudo-scalar. If we denote the Lorentz boost for a Dirac spinor as $D(a)$, we can relate the Floquet-Bloch solution in the lab frame |$\ket{\Psi_{\alpha\bm{k}}(x)}$ to that in the rest frame |$\ket{\Psi^{\prime}(x^{\prime})}$ as,
|$\ket{\Psi_{\alpha\bm{k}}(x)}=D(a^{-1})\ket{\Psi^{\prime}_{\alpha\bm{k}^{\prime}}(ax)}=\frac{1}{\sqrt{N}}e^{-i(a^{-1}k^{\prime})\cdot x}D(a^{-1})\ket{\Phi^{\prime}_{\alpha\bm{k}^{\prime}}(q\cdot x)}.$ (27)
Note that |$\ket{\Psi^{\prime}(x^{\prime})}$ is to be determined by solving the Floquet-Bloch equation Eqn. (8) using the parameters $q^{\prime},V^{\prime},A^{\prime}_{5}{}^{\mu},\ldots$ in the rest frame. From Eqn. (27), we find that the quasi-momentum-energy vector and the Floquet-Bloch state in the two frames are related by $k=a^{-1}k^{\prime}$ and |$\ket{\Phi(q\cdot x)}=D(a^{-1})\ket{\Phi^{\prime}(q\cdot x)}$. The former relation implies that the spectra in the two frames are related by Lorentz boost (See Fig. 2 below for an example).
As a demonstration, let us consider a propagating wave defined by
$V=\frac{A}{2}\left(\Gamma^{y}-i\Gamma^{z}\right)$ (28)
with $A$ determining its strength. Physically, this wave represents circularly polarized laser fields propagating in the $x$-direction with $A=E/\omega$ related to the electric field $E$. This wave is invariant under Lorentz boost in the $x$-direction and thus $V^{\prime}(x^{\prime})=V(ax)$. Below, we assume that we can select its speed $v$ freely.
Figure 2: Floquet quasi-energy spectra of massless Dirac electrons in a circularly polarized light wave Eqn. (28) for the (a) time-like ($\omega=10,\;Q=5$) and (b) space-like ($\omega=5,Q=10$) cases. The left figures show the spectra plotted against $k_{x}$ and $k_{y}$ while fixing to $k_{z}=0$, while the middle figures are their $k_{y}=k_{z}=0$ slices. The right figures show the corresponding spectra in the rest frame with the Lorentz contracted parameters Eqn. (23), (24). The spectra in the lab and rest frames are related by Lorentz boost $k=a^{-1}k^{\prime}$, which tilts the spectra. The green dots are the Weyl point that are pair created from the Dirac point at $\varepsilon=0,\bm{k}=\bm{0}$ due to the Floquet chiral gauge fields. The green dots in the time-like case are the emergent Floquet Weyl nodes appearing by resonance of photo-dressed states Bucciantini et al. (2017); Hirai et al. (2024). Blue shaded areas in the right figures are the first Brillouin zones, i.e., (a) $\varepsilon\in\left[-\Omega/2,\Omega/2\right]$ and (b) $k_{x}\in\left[-Q/2,Q/2\right]$, in the rest frame, while those in the middle figures are the first Brillouin zones in the lab frame.
Massless Dirac electron and emergent Weyl points.— First, we consider the massless case $m=0$ with $A_{\mu}=0$ and $A^{5}=0$. Figure 2 depicts typical Floquet-Bloch spectra for the electrons in the (a) time like and (b) space like waves. By comparing the spectra in the lab and rest frames, we see that they are related by the Lorentz boost $k=a^{-1}k^{\prime}$ as expected. The blue shaded regions are the 1st Brillouin zones. We also notice that Weyl points appear that did not exist in the undriven case. The green dots in the figures correspond to Weyl points that are pair-created from the original Dirac point $(\varepsilon=0,\bm{k}=0)$ due to the effective chiral gauge field  Wang et al. (2014); Ebihara et al. (2016); Chan et al. (2016); Hübener et al. (2017); Yoshikawa et al. (2022). The orange dots corresponds to emergent Weyl points that are created by the hybridization of photo-dressed bands Bucciantini et al. (2017); Hirai et al. (2024).
Effective chiral gauge field is generated in Dirac and Weyl semimetals when circularly polarized laser fields is applied Ebihara et al. (2016). This can be described by the effective Floquet Hamiltonian within the high frequency expansion $H_{\rm eff}=H_{\bf k}+\frac{1}{\omega}[V,V^{\dagger}]+\mathcal{O}(1/\omega^{2})$ Kitagawa et al. (2011), which gives the chiral gauge field in the rest frame to the lowest order in $1/\omega$ as
$\bm{A}_{5{\rm eff}}^{\prime}{}=\left(\gamma\frac{v_{F}^{2}A^{2}}{\omega},0,0\right).$ (29)
We can use the Lorentz transformation Eqn. (25) to obtain the effective chiral gauge field in the lab frame
$A_{5{\rm eff}}^{\mu}=\left(\frac{v_{F}/v}{1-v_{F}^{2}/v^{2}}\frac{v_{F}^{2}A^{2}}{\omega},\frac{1}{1-v_{F}^{2}/v^{2}}\frac{v_{F}^{2}A^{2}}{\omega},0,0\right).$ (30)
This predicts that a the 00-th component of the chiral gauge field, also known as the chiral chemical potential $\mu_{5}$, is generated. In principle this can be detected by chiral magnetic effect under the application of magnetic field $J=\frac{e}{h}\mu_{5}B$ Fukushima et al. (2008). An important physical consequence is that the chiral gauge field is enhanced by Lorentz contraction $\gamma$. If one can realize propagating wave with speed approaching the Fermi velocity, the Floquet engineering effect will become largely enhanced.
Figure 3: (a) Phase diagram and spectra of the massive Dirac fermion with mass $m=1$ and field strength $A=4$ driven by a circularly polarized light wave Eqn. (28). The color corresponds to size of the gap at $\bm{k}=0$ (b) The Floquet-Bloch spectrum at $\bm{k}=0$ plotted against frequency $\omega$ with fixed $Q=10$. (i-iv) The Floquet-Bloch spectra for several values of frequency with fixed $Q=10$. The green dots in (ii), (iv) are the 3D Weyl nodes. The red color in the figures  (i)-(iv) is the static weight of the Floquet states defined by $I_{\alpha\bm{k}}=|\ket{\phi^{0}_{\alpha\bm{k}}}|^{2}$ which roughly indicates how strong the state will couple to static local electron baths.
Massive Dirac electron and phase transitions.— Next, let us study the case of a massive Dirac electron in the propagating wave Eqn. (28). We focus on the low energy Floquet-Bloch bands that descend from the upper and lower Dirac bands. In the time-like case, we expect the spectrum at large $\omega$ to be given by
$\varepsilon^{\prime}_{\alpha}(\bm{k}^{\prime})=\pm v_{F}\sqrt{(\sqrt{k_{x}^{\prime}+m^{2}}\pm A_{5{\rm eff}}^{\prime x})^{2}+k_{y}^{\prime 2}+k_{z}^{\prime 2}}$ (31)
in the rest frame Ebihara et al. (2016). The gap closes when the strength of the chiral gauge field exceeds the mass parameter $A_{5{\rm eff}}^{\prime x}=m$. In Fig. 3 (i)-(iv), we depict the Floquet-Bloch spectrum for several values of $\omega$ while fixing the wave number $Q=10$. In the case of large $\omega$ as in Fig. 3 (i), the upper and lower Dirac bands split by the chiral gauge field and the gap at $\bm{k}=0$ becomes small. A phase transition occur to a state with two Weyl fermions as depicted in Fig. 3 (ii) when the gap close at $A_{5{\rm eff}}^{\prime x}=m$. Near the phase transition, the Weyl fermions are type-II Soluyanov et al. (2015) since the bands are tilted by the Lorentz boost. When we further approach the light-like limit $\omega\sim Q$, the energy spectrum becomes dense since the Brillouin zone becomes narrow and multiple band foldings occur. This is the region where the system approaches the adiabatic limit $\omega^{\prime}\to 0$ in the rest frame. In this limit, we expect adiabatic time evolution with nonperturbative geometric effects dominates the dynamics Takayoshi et al. (2021b). In the space like case, the gap is always open, while the bands are tilted and shows a chiral flow (Fig. 3 (iv)). Weyl points can exist if $A_{5{\rm eff}}^{\prime 0}>m$ is satisfied in the rest frame.
Discussion.— We have studied the Floquet-Bloch band engineering of Dirac fermions induced by propagating waves. The Lorentz covariance of the Dirac equation allows us to classify the waves into three categories: time-like, light-like, and space-like, and the corresponding electronic bands are described by Floquet states, Volkov states Volkov (1935), and Bloch states, respectively. One of the most striking findings is that the Floquet-Bloch bands undergo significant modifications near the light-like case due to the Lorentz contraction factor $\gamma$ appearing in the effective Floquet terms as in Eqns. (29, 30). The ability to design propagating waves with specific characteristics could enable efficient realization of Floquet engineering of band topology in quantum materials.
ACKNOWLEDGMENTS
The authors appreciate the fruitful discussions with S. Kitamura, S. Chaudhary, and K. Saha. This work is supported by JSPS KAKENHI (No. JP23H04865, No. JP23K25837 and No. JP23K22487), MEXT, Japan, and JST CREST Grant No. JPMJCR19T3, Japan.